# ]https://zf-wei.github.io

### v2

@@ I Introduction
Community detection using graph embeddings relies on representing graph nodes as points in a low-dimensional vector space, where communities manifest as clusters of points that are close to each other and sufficiently separated from other clusters [14, 15, 16]. These clusters are then recovered using standard data clustering techniques such as $k$-means clustering [17]. Given the practical importance of consistently detecting communities and the effectiveness of graph embedding approaches in doing so, it is essential to investigate the robustness of graph embeddings for community detection under network perturbations. In this context, robustness refers to the ability of a graph embedding method to tolerate perturbations while still accurately recovering the original community structure [18, 19, 20]. This investigation is crucial for understanding how communities identified by graph embedding methods withstand random errors or adversarial attacks, serving as a proxy for evaluating network functionality. Robust graph embeddings ensure the reliability and stability of community detection algorithms in real-world network applications.
@@
Previous research has investigated the impact of network perturbations on the robustness of community structures, focusing on techniques that characterize structural properties of communities such as internal link density [21, 20]. However, little attention has been given to understanding the robustness of graph embedding methods used for community detection under perturbations in the graph topology. While there are numerous graph embedding techniques available for community detection [1, 22], it remains challenging for practitioners to select a robust method and comprehend its implications across various factors, including the size of the network, strength of the original community structure, type of perturbation, and the specific embedding technique used. Moreover, the majority of research on using graph embeddings for community detection has been conducted on synthetic networks, warranting validation of the effects of perturbations on real-world network data. Addressing these gaps in research, we aim to investigate the robustness of graph embedding methods for community detection under different types of perturbations in the graph topology. By exploring these factors across synthetic and real-world network datasets, we seek to provide insights into the effectiveness and limitations of graph embedding techniques in preserving community structures amidst network perturbations.
@@
In this paper, we address this gap of knowledge by devising a systematic framework to investigate the robustness of graph embedding methods used to extract community structures under network perturbations. We consider both synthetic Lancichinetti-Fortunato-Radicchi (LFR) benchmark graphs [23] and empirical networks with degree and community size heterogeneity. Our perturbation approach involves the removal of edges based on random node selection [21] and targeted node selection using betweeness centrality [24, 25, 26]. We use seven state-of-the-art graph embedding methods, namely Laplacian eigenmap (LE) [27], locally linear embedding (LLE) [28], higher-order preserving embedding (HOPE) [29], modularized non-negative matrix factorization (M-NMF) [30] in the matrix factorization family; and DeepWalk [31], large-scale information network (LINE) [32], and node2vec [33] in the random walk family. We use $k$-means based on spherical distance [34] to achieve more effective clustering results in high-dimensional spaces. We use element-centric similarity (ECS) [35], a widely adopted community similarity metric, to quantify changes in the similarity partitions after many realizations with perturbations as a proxy of characterizing graph embedding robustness. This comprehensive methodology allows us to systematically assess and compare the impact of various perturbation strategies on the robustness of graph embedding methods for community detection in both synthetic and real-world networks.
@@
Our findings indicate that incremental selection of nodes and subsequent removal of their adjacent edges consistently reduce ECS. However, this impact tends to be more noticeable across graph embedding methods, but node2vec and LLE, and varies across network sizes, the initial community structure strength, and the perturbation type. This aligns with the expectation that removing a greater number of edges results in a more substantial disruption of the community structure. Notably, targeted node selection leads to a more pronounced and accelerated decline in ECS compared to random node selection. This is rationalized by the fact that adversarial attacks tend to swiftly dismantle the network’s community structure. Among the seven network embedding methods considered, node2vec and LLE consistently prove to be more robust for community detection within their respective graph embedding families; these two methods provide the highest ECS similarity scores. Overall, node2vec is the most robust graph embedding method for community detection. Our results generally concur with those of Kojaku et al. [16], which found that node2vec learns more consistent community structures with networks of heterogeneous degree distributions and size.
@@
We have made available the code to reproduce all the results at  [^1] https://github.com/zf-wei/Robustness-of-Graph-Embeddings-for-Community-Detection (accessed on April 29, 2024).
II Methods
II.1 LFR benchmark
@@
Lancichinetti–Fortunato–Radicchi (LFR) benchmark graphs [23, 37] have become a widely adopted tool for generating graphs with ground-truth communities. One of the primary advantages of utilizing the LFR benchmark is its ability to produce graphs where both degree and community size distributions follow power-law distributions. These distributions closely resemble the properties observed in many real-world networks [23, 38, 39, 40]. The exponents governing the degree and community size distributions are controlled by the parameters $\alpha$ and $\beta$, respectively. To generate LFR benchmark graphs, one must specify various other parameters, including the average degree $\langle k\rangle$, maximum degree $maxk$, minimum community size $minc$, maximum community size $maxc$, and the mixing parameter $\mu$. The mixing parameter $\mu$ represents the fraction of nodes that share edges across different communities, with lower values of $\mu$ leading to a higher ratio between internal and external edges. This results in increased modularity, a commonly used quality metric for assessing community strength in community research [11]. Thus, smaller values of $\mu$ correspond to stronger partitions. Note that LFR benchmark graphs are undirected and unweighted.
@@
We generate LFR benchmark graphs using the LFR_benchmark_graph method in the publicly available Python package NetworkX [41]. Our experiments involve two different network sizes: $1,000$ nodes and $10,000$ nodes, allowing us to investigate the impact of perturbations at varying scales. Furthermore, we explore the influence of the initial community strength by adjusting the mixing parameter, $\mu$. The specific parameter values utilized for the generation of LFR benchmark graphs are detailed in Table 1. Note that the proposed experimental design can be applied to any other synthetic graph generators such as the Girvan–Newman benchmark [42] and the stochastic block model [43].
@@
$N$ number of nodes $1,000$ and $10,000$
$maxk$ max degree for LFR $0.1N$
$\langle k\rangle$ average degree for LFR $25$
$maxc$ max community size for LFR $0.1N$
$\alpha$ degree distribution exponent for LFR $-2$
$\beta$ community size distribution exponent for LFR $-1.1$
$\mu$ mixing parameter for LFR $0.01,0.1,0.2,0.3,0.4,0.5$
$r$ number of realizations $50$
II.2 Edge removal procedure
@@
Let’s first discuss edge removal based on random node selection. This simulates random network errors in real-world scenarios [44, 45]. We do so by selecting $p\%$ of nodes randomly and deleting their adjacent edges. To make our experiment unbiased, we independently generate $50$ distinct sets of nodes. In other words, the number of realizations is $r=50$ (see Table 1). We ensure that the removal of edges from each set of nodes does not disconnect the remaining part of the network, which is required by the graph embedding methods. In the event that the remaining part of the network becomes disconnected after the removal of a set of edges, we omit this set of nodes and proceed with re-sampling. After each removal, we measure the similarity between the community structure of the remaining network in the embedding space calculated using data clustering (i.e., $k$-means; see more details in Sec. II.3) and the community structure of the original LFR network using community similarity metrics (i.e., ECS [35]) as detailed in Sec. II.4. The overall similarity score for a given $p\%$ value is calculated by averaging the similarity scores across all $50$ sets.
@@
Second, we discuss targeted edge removal. Targeted edge removal is based on the betweenness centrality of nodes. Betweenness centrality measures the importance of a node in a network based on how often it lies on the shortest paths between other nodes [46]. Nodes with high betweenness centrality usually indicate nodes of vital importance to the network, so removing edges adjacent to these nodes might cripple the functionality of the network [47]. This approach models an adversarial attack in real-world situations [48, 49]. To implement edge removal based on targeted node selection, we first rank the nodes in the network by their betweenness centrality. The node with the lowest betweenness centrality has rank $1$, and the node with the highest betweenness centrality has rank $N$, where $N$ is the number of nodes in the network. We then independently generate $50$ sets of nodes, each containing $p\%$ of nodes in the network. The nodes in each set are selected by randomly sampling nodes with probability proportional to their betweenness centrality rank. We ensure that the removal of edges adjacent to each set of nodes does not disconnect the remaining part of the network, which is required by graph embedding methods. In the event that the remaining part of the network becomes disconnected after the removal of a set of edges, we omit the set of nodes and proceed with re-sampling. After each removal, we measure the similarity between the community structure of remaining network in the embedding space computed through data clustering (i.e., $k$-means; see more details in Sec. II.3) and the ground-truth community structure of the original LFR network using ECS. The overall similarity score for a given $p\%$ is calculated by averaging community similarity scores across all $50$ sets.
In our experiments, the proportion of selected nodes starts from $5\%$ with an increment of $5\%$. The upper bound of proportion depends on the network structure and the node selection method. For example, when we select a network with $N=1,000$, $\mu=0.01$, and use edge removal based on targeted node selection, it is difficult to keep the network connected when deleting edges adjacent to $35\%$ of nodes. In these cases, we use a proportion of selected nodes that is below the threshold, e.g., $5\%$, $10\%$, $15\%$, $20\%$, $25\%$, and $30\%$. In addition, we also perform data clustering on the embeddings of the original LFR networks with no edges removed and compare the results with the ground-truth community structure. This result is recorded as “0% of nodes selected.” We also capture the variation of the measurements by computing standard deviations. As the standard deviations are negligible with respect to the mean, we do not show those results here.
II.3 Data clustering
@@
II.4 Community similarity
To measure the similarity between community partitions, we employ ECS [35], computed with the CluSim package [61]. Compared to other clustering similarity metrics such as the normalized mutual information (NMI) [62], ECS is a most robust metric for comparing clusters and better addresses challenges such as biases in randomized membership, skewed cluster sizes, and the problem of matching [35]. We use default parameter values in ECS computation. ECS yields values in the range from 00 to $1$, where a higher value indicates a greater similarity between partitions.

### v3

@@ I Introduction
Community detection using graph embeddings relies on representing graph nodes as points in a low-dimensional vector space, where communities manifest as clusters of points that are close to each other and sufficiently separated from other clusters [14, 15, 16]. These clusters are then recovered using standard data clustering techniques such as $k$-means clustering [17]. Given the practical importance of consistently detecting communities and the effectiveness of graph embedding approaches in doing so, it is essential to investigate the robustness of graph embeddings for community detection under network perturbations. In this context, robustness refers to the ability of a graph embedding method to tolerate perturbations while still accurately recovering the original community structure [18, 19, 20]. This investigation is crucial for understanding how communities identified by graph embedding methods withstand random errors or adversarial attacks, serving as a proxy for evaluating network functionality. Robust graph embeddings ensure the reliability and stability of community detection algorithms in real-world network applications. Moreover, robustness in graph embeddings has practical implications beyond community detection. Robustness to edge deletion in community detection reflects a method’s ability to recover meaningful structure despite incomplete or perturbed network data. In recommendation systems, such robustness ensures that user or item communities remain stable even in the presence of missing interactions, thereby supporting consistent and reliable recommendations. In anomaly detection, it enhances the capacity to distinguish between genuine structural anomalies and incidental perturbations in network communication systems, improving anomaly detection accuracy.
@@
Previous research has investigated the impact of network perturbations on the robustness of community structures using traditional community detection methods [21, 20]. However, little attention has been given to understanding the robustness of graph embedding methods used for community detection under perturbations in the graph topology. While there are numerous graph embedding techniques available for community detection [1, 22], it remains challenging for practitioners to select a robust method and comprehend its implications across various factors, including the size of the network, strength of the original community structure, type of perturbation, and the specific embedding technique used. Moreover, the majority of research on using graph embeddings for community detection has been conducted on synthetic networks, warranting validation of the effects of perturbations on real-world network data. Addressing these gaps in research, we aim to evaluate the robustness of graph embedding methods for community detection under various types of graph topology perturbations, and to identify the methods that demonstrate the highest robustness. By exploring these factors across synthetic and real-world network datasets, we seek to provide insights into the effectiveness and limitations of graph embedding techniques in preserving community structures amidst network perturbations.
In this paper, we address this gap of knowledge by devising a systematic framework to investigate the robustness of graph embedding methods used to extract community structures under network perturbations. We consider both synthetic Lancichinetti-Fortunato-Radicchi (LFR) benchmark graphs [23] and empirical networks with degree and community size heterogeneity.
@@
Our perturbation approach involves the removal of edges based on random node selection [21] and targeted node selection using betweeness centrality [24, 25, 26]. We use seven state-of-the-art graph embedding methods, namely Laplacian eigenmap (LE) [27], locally linear embedding (LLE) [28], higher-order preserving embedding (HOPE) [29], modularized non-negative matrix factorization (M-NMF) [30] in the matrix factorization family; and DeepWalk [31], large-scale information network (LINE) [32], and node2vec [33] in the random walk family. We use $k$-means based on spherical distance [34] to achieve more effective clustering results in high-dimensional spaces. We use element-centric similarity (ECS) [35], a widely adopted community similarity metric, to quantify changes in the community structure after perturbations as a proxy of characterizing graph embedding robustness. This comprehensive methodology allows us to systematically assess and compare the impact of various perturbation strategies on the robustness of graph embedding methods for community detection in both synthetic and real-world networks.
@@
Our findings indicate that incremental selection of nodes and subsequent removal of their adjacent edges reduce ECS. However, this impact tends to be more noticeable across graph embedding methods, but node2vec and LLE, and varies across network sizes, the initial community structure strength, and the perturbation type. This aligns with the expectation that removing a greater number of edges results in a more substantial disruption of the community structure. Notably, targeted node selection leads to a more pronounced and accelerated decline in ECS compared to random node selection. This is rationalized by the fact that adversarial attacks tend to swiftly dismantle the network’s community structure. Among the seven network embedding methods considered, node2vec and LLE consistently prove to be more robust for community detection within their respective graph embedding families; these two methods provide the highest ECS similarity scores. Overall, node2vec is the most robust graph embedding method for community detection. Our results generally concur with those of Kojaku et al. [16], which found that node2vec learns more consistent community structures with networks of heterogeneous degree distributions and size.
We notice that community detection methods based on graph neural networks (GNNs) currently constitute a popular approach [36]. The focus of our study, however, is on shallow graph embedding methods rather than GNNs.
@@
We have made available the code and data to reproduce all the results at  [^1] https://t.ly/uIclh (accessed on August 1, 2025),  [^2] https://t.ly/OzFzV (accessed on August 1, 2025).
II Related Work
II.1 Prior Work Closely Related to the Present Study
There has been substantial research in network community detection leveraging graph embedding techniques.
Tandon et al. [14] tested the ability of the aforementioned seven state-of-the-art graph embedding techniques to detect communities on benchmark graphs. It was reported that if the parameters of the embedding techniques are suitably chosen, the performance of community detection methods based on graph embedding techniques is comparable to the traditional community detection algorithms, including Infomap and Louvain.
Zhang et al. [39] proposed a method that simultaneously performs network embedding and community detection in directed networks. Their model introduced separate vector representations for out-nodes and in-nodes, allowing nodes to belong to different out- and in-communities. Their method has improved performance in community detection while jointly estimating node embeddings and their community assignments. An efficient alternating optimization scheme was developed, and theoretical guarantees for both embedding and community detection are provided, supported by experiments on simulated and real-world networks.
Ghanbari et al. [40] investigated the efficiency of different node embedding algorithms combined with clustering methods for community detection. They evaluated four embedding techniques, including leveraging convolution, attention, inductivity, and shallowness, paired with three clustering algorithms, including KMeans++, DBSCAN, and hierarchical clustering. Experimental results showed that GraphSAGE, an inductive embedding method, combined with KMeans++ produced the best performance.
Zhu et al. [41] proposed a structural equivalence embedding method based on non-negative matrix factorization (SENMF) for community detection, which embeds two similarity measures one capturing node closeness and the other reflecting similarity between distant nodes into a low-dimensional space. Experimental results show that SENMF outperforms several well-known network embedding and traditional community detection methods. In fact, as mentioned in [41], when taking parameters $\theta=0$ and $\varphi=0$, SENMF will equal M-NMF.
Kovács et al. [42] proposed an iterative embedding and reweighting procedure that strengthens intra-community links while weakening inter-community links, enhancing cluster visibility and detectability. This process can lead to such strong geometric separation that even simple link parsing recovers communities with high precision. Additionally, it can serve as a pre-processing step to improve traditional community detection algorithms, including Louvain, Infomap, and HDBSCAN.
Pankratz et al. [43] proposed a method to enhance community detection by initializing heuristic algorithms with partitions obtained from network embeddings. Their approach embeds nodes into a latent space to filter noise while preserving structural relationships, followed by clustering to generate a stable partition. This initialization improves the performance and reduces the volatility of traditional methods such as Louvain and Leiden, particularly in networks with weak community structures. Experimental results demonstrated that the effectiveness of this procedure depends on network topology rather than size.
Zhu et al. [44] proposed a novel Graph Convolutional Network (GCN) method for unsupervised community detection, termed UCDMI. This method leverages mutual information maximization to capture fine-grained information about the global network structure without supervision. A new aggregation function for GCN is introduced to differentiate the importance of neighboring nodes, enhancing the quality of node representations. The proposed approach demonstrates significant improvements in community detection performance through extensive experiments, outperforming several state-of-the-art methods. This method offers a promising direction for unsupervised community detection by integrating mutual information maximization with GCNs.
In this work, we focus on community detection in undirected networks using the seven graph embedding techniques from [14]. Rather than introducing new community detection methods, we examine the robustness of these embedding techniques under network perturbations.
II.2 Other Prior Work Related to the Present Study
The following articles examines the robustness of community structures using traditional community detection methods under different types of network perturbations.
Karrer et al. [18] proposed a framework for evaluating the significance of community structures by perturbing networks with configuration model random graphs and measuring the resulting changes using variation of information. They employed a spectral optimization method for modularity to perform community detection. This approach was applied to a wide range of networks, including both real-world and synthetic examples.
Ramirez-Marquez et al. [45] analyzed the effects of weight uncertainties on community structures in networks with weighted edges. Using Monte Carlo-based uncertainty propagation, they analyzed how variations in link weights influence community structures. The approach was demonstrated through the analysis of an electric power system network, highlighting the significance of weight uncertainties in determining reliable community partitions.
Peche et al. [46] investigated perturbations where the connections between network nodes are influenced by a random geometric graph. They demonstrated that community detection using spectral methods remains robust to such noise, based on a thorough and detailed analysis of the spectrum of the random graph’s adjacency matrix.
Jia et al. [47] proposed a robustness guarantee for community detection against adversarial structural perturbations. For any given community detection method, they introduced a smoothed community detection approach by randomly perturbing the graph structure. They theoretically demonstrated that the smoothed community detection method provably groups a given arbitrary set of nodes into the same community (or different communities) provided that the number of edges added or removed by an attacker is bounded. The method was also empirically evaluated on multiple real-world graphs with ground truth communities.
Tian et al. [20] studied the robustness of community detection algorithms under edge addition perturbations in synthetic and empirical temporal networks. They considered two edge addition scenarios—random and targeted—and evaluated the robustness of communities detected by Infomap, Label Propagation, Leiden, and Louvain algorithms. It was found that robustness performance varies across different community similarity metrics and detection methods, highlighting the dependence of community stability on the choice of algorithm.
Furthermore, given that graph neural network-based community detection methods are numerous and currently constitute a popular approach, to reflect the current research trends, we include the following works on graph community detection using graph neural networks.
Zhang et al. [48] proposed a deep learning-based approach for dynamic community detection that incorporates node relevancy to capture structural information. Additionally, it employs cross-entropy to smooth learned features between adjacent time steps, addressing potential large discrepancies that don’t align with real-world dynamics.
Sun et al. [49] introduced a novel graph neural network encoding method for multiobjective evolutionary algorithms to address community detection in complex attribute networks. The approach transforms the problem into one with continuous decision variables, facilitating smoother fitness landscapes, and demonstrates superior performance on both single- and multi-attribute networks.
[…]
Su et al. [36] provided a comprehensive survey of recent advancements in deep learning techniques for community detection, particularly highlighting their utility in handling high-dimensional network data. It proposes a new taxonomy that categorizes state-of-the-art methods, including various deep neural network models, and summarizes benchmark datasets, evaluation metrics, and open-source implementations.
To address the limitations of previous research, such as the lack of a systematic evaluation of edge removal’s impact on community structure robustness, our study examines edge deletion based on both random and targeted node selection. We apply this approach to both synthetic and real-world networks, accounting for degree and community size heterogeneity. This perturbation approach presents a novel method of network perturbation with real-world relevance. To compare the community structures before and after the network perturbation, we utilized ECS, a robust metric for evaluating network clusters.
III Methods
III.1 LFR Benchmark
@@
Lancichinetti–Fortunato–Radicchi (LFR) benchmark graphs [23, 51] have become a widely adopted tool for generating graphs with ground-truth communities. One of the primary advantages of utilizing the LFR benchmark is its ability to produce graphs where both degree and community size distributions follow power-law distributions. These distributions closely resemble the properties observed in many real-world networks [23, 52, 53, 54]. The exponents governing the degree and community size distributions are controlled by the parameters $\alpha$ and $\beta$, respectively. To generate LFR benchmark graphs, one must specify various other parameters, including the average degree $\langle k\rangle$, maximum degree $maxk$, maximum community size $maxc$, and the mixing parameter $\mu$. The mixing parameter $\mu$ represents the fraction of nodes that share edges across different communities, with lower values of $\mu$ leading to a higher ratio between internal and external edges. This results in increased modularity, a commonly used quality metric for assessing community strength in community research [11]. Thus, smaller values of $\mu$ correspond to stronger partitions. Note that LFR benchmark graphs are undirected and unweighted.
@@
We generate LFR benchmark graphs using the LFR_benchmark_graph method in the publicly available Python package NetworkX [55]. Our experiments involve two different network sizes: 1,0001,0001 , 000 nodes and 10,00010,00010 , 000 nodes, allowing us to investigate the impact of perturbations at varying scales. Furthermore, we explore the influence of the initial community strength by adjusting the mixing parameter, $\mu$. The specific parameter values utilized for the generation of LFR benchmark graphs are detailed in Table 1. Note that the proposed experimental design can be applied to any other synthetic graph generators such as the Girvan–Newman benchmark [56] and the stochastic block model [57].
@@
$N$ number of nodes 1,0001,0001 , 000 and 10,00010,00010 , 000
[…]
$\langle k\rangle$ average degree for LFR 252525
[…]
$\alpha$ degree distribution exponent for LFR 2-2- 2
$\beta$ community size distribution exponent for LFR 1.1-1.1- 1.1
$\mu$ mixing parameter for LFR 0.01,0.1,0.2,0.3,0.4,0.50.01,0.1,0.2,0.3,0.4,0.50.01 , 0.1 , 0.2 , 0.3 , 0.4 , 0.5
$r$ number of realizations 505050
III.2 Edge Removal Procedure
@@
Let’s first discuss edge removal based on random node selection. This simulates random network errors in real-world scenarios [58, 59]. We do so by selecting $p\%$ of nodes randomly and deleting their adjacent edges. To make our experiment unbiased, we independently generate 505050 distinct sets of nodes. In other words, the number of realizations is $r=50$ (see Table 1). We ensure that the removal of edges from each set of nodes does not disconnect the remaining part of the network, which is required by the graph embedding methods. In the event that the remaining part of the network becomes disconnected after the removal of a set of edges, we omit this set of nodes and proceed with re-sampling. After each removal, we measure the similarity between the community structure of the remaining network calculated using data clustering in the embedding space (i.e., $k$-means; see more details in Sec. III.3) and the community structure of the original LFR network using community similarity metrics (i.e., ECS [35]) as detailed in Sec. III.4. The overall similarity score for a given $p\%$ value is calculated by averaging the similarity scores across all 505050 sets.
@@
Second, we discuss targeted edge removal. Targeted edge removal is based on the betweenness centrality of nodes. Betweenness centrality measures the importance of a node in a network based on how often it lies on the shortest paths between other nodes [60]. Nodes with high betweenness centrality usually indicate nodes of vital importance to the network, so removing edges adjacent to these nodes might cripple the functionality of the network [61]. This approach models an adversarial attack in real-world situations [62, 63]. To implement edge removal based on targeted node selection, we first rank the nodes in the network by their betweenness centrality. The node with the lowest betweenness centrality has rank 111, and the node with the highest betweenness centrality has rank $N$, where $N$ is the number of nodes in the network. We then independently generate 505050 sets of nodes, each containing $p\%$ of nodes in the network. The nodes in each set are selected by randomly sampling nodes with probability proportional to their betweenness centrality rank. We ensure that the removal of edges adjacent to each set of nodes does not disconnect the remaining part of the network, which is required by graph embedding methods. In the event that the remaining part of the network becomes disconnected after the removal of a set of edges, we omit the set of nodes and proceed with re-sampling. After each removal, we measure the similarity between the community structure of remaining network computed through data clustering in the embedding space (i.e., $k$-means; see more details in Sec. III.3) and the ground-truth community structure of the original LFR network using ECS. The overall similarity score for a given $p\%$ is calculated by averaging community similarity scores across all 505050 sets.
[…]
III.3 Data Clustering
@@
III.4 Community Similarity
To measure the similarity between community partitions, we employ ECS [35], computed with the CluSim package [75]. Compared to other clustering similarity metrics such as the normalized mutual information (NMI) [76], ECS is a most robust metric for comparing clusters and better addresses challenges such as biases in randomized membership, skewed cluster sizes, and the problem of matching [35]. We use default parameter values in ECS computation. ECS yields values in the range from 0 to 111, where a higher value indicates a greater similarity between partitions.